# DynamicFace: High-Quality and Consistent Video Face Swapping using Composable 3D Facial Priors

### v1

@@ Abstract
Face swapping transfers the identity of a source face to a target face while retaining the attributes like expression, pose, hair, and background of the target face. Advanced face swapping methods have achieved attractive results. However, these methods often inadvertently transfer identity information from the target face, compromising expression-related details and accurate identity. We propose a novel method DynamicFace that leverages the power of diffusion model and plug-and-play temporal layers for video face swapping. First, we introduce four fine-grained facial conditions using 3D facial priors. All conditions are designed to be disentangled from each other for precise and unique control. Then, we adopt Face Former and ReferenceNet for high-level and detailed identity injection. Through experiments on the FF++ dataset, we demonstrate that our method achieves state-of-the-art results in face swapping, showcasing superior image quality, identity preservation, and expression accuracy. Besides, our method could be easily transferred to video domain with temporal attention layer. Our code and results will be available on the project page: https://dynamic-face.github.io/.
Introduction
@@
Recent works (Li et al. 2020; DeepFakes 2019; Shen, Yan, and Zeng 2018) have made great efforts to achieve good face swapping results. However, these methods often focus on inner facial texture but ignore shape and illumination. Later, some works (Wang et al. 2021; Xu et al. 2021) combine the expert knowledge from 3D face reconstruction to achieve better performance with both reconstruction loss and identity loss. But training model on two losses will face a trade-off problem between identity and reconstruction, which often leads to instability and unreliability. Balancing identity and reconstruction with only losses constraint is conflict. We should provide accurate information for controllable generation.
Recently, diffusion-based models (Ho, Jain, and Abbeel 2020; Song et al. 2021; Rombach et al. 2022) have exhibited high customizability for various conditions and impressive ability to generate images with high resolution and complex scenes. Some works (Han et al. 2024; Zhao et al. 2023; Kim et al. 2022) try to utilize the diversity and powerful generation ability of diffusion model to achieve face swapping. However, these methods could not achieve video face swapping directly. And idealized face guidance should preserve the non-identity attributes of $I_{tgt}$, which include non-facial attributes (e.g., background and hair), facial posture (e.g., expression and pose) and facial color (e.g., lighting). Previous works didn’t pay much attention to the potentiality of 3D facial priors to simultaneously maintain identity from source face and complicated motion information from target face.
@@
In this work, we propose a method named DynamicFace for face swapping by incorporating precise and disentangled facial conditions to the powerful Stable Diffusion model (Rombach et al. 2022) and apply temporal attention layer in AnimateDiff (Guo et al. 2024) to make DynamicFace also available in video domain. We introduce a new multiple conditioning mechanism to address the limitations of previous methods. Specifically, 1) to accurately control expression, pose, and other motion, we introduce four fine-grained facial conditions: background, shape-aware normal map, expression-related landmark, and identity-removed UV texture map. These four conditions fully encapsulate the desired information in the target face and are well disentangled from each other. 2) to avoid introducing unexpected identity information from the target face, we propose a alignment strategy to form spatial-aligned conditions using 3D facial priors. This involves extracting pose and expression parameters from the target face video and shape parameters from the source image using the prior of 3D face reconstruction model 3DDFA-V3 (Wang et al. 2024b), and then rendering the 2D face normal map and the UV texture map with these facial parameters.
In summary, our main contributions are as follows:
@@
  • We adopt an identity-aware ReferenceNet and Face Former to inject facial information. These two modules ensure the identity consistency of the human subject across kinds of expressions and poses without any identity loss.
Related Work
@@
Early methods directly warp the source face according to the target facial landmarks, thus failing to address the large differences in posture and expression. 3DMM-based methods swap faces by 3D-fitting and re-rendering. However, these methods often cannot handle lighting differences on face and suffer from poor fidelity. Later, GAN-based methods improve the fidelity of the generated faces. Deepfakes (DeepFakes 2019) transfers the target attributes to the source face by an encoder-decoder structure while being constrained by two specific identities. FSGAN (Nirkin, Keller, and Hassner 2019) animates the source face with target facial landmarks and fuses it with the target background. SimSwap (Chen et al. 2020) introduces a feature matching loss to help preserve the target attributes. These methods require a facial mask for blending, which limits face shape variation.
@@
Diffusion models have shown remarkable performance in various generative tasks, emerging as a powerful alternative to traditional GANs for image synthesis. The innovative approach of these models lies in their ability to gradually refine an image by reversing a learned noising process, which has led to significant advancements in image quality and diversity. In the context of face swapping, diffusion models have introduced a new paradigm for identity manipulation in face images. DiffSwap (Zhao et al. 2023) proposes a diffusion-based pipeline in latent space for face swapping by integrating 3D-aware masked diffusion and midpoint estimation during the reverse process. DiffFace (Kim et al. 2022) introduces an ID conditional DDPM. However, pure pixel-wise diffusion without pretraining requires more data, and extra forward steps with gradients make training more memory-consuming. These both make training process difficult. Face Adapter (Han et al. 2024) only utilizes an efficient adapter to introduce identity information and preserve attributes of the target face image, but it is difficult to achieve temporal consistency when balancing the trade-off between identity and motion. In addition, these diffusion-based methods could not easily transfer to the video domain.
@@
Figure 3: The overview of the proposed method. A VAE encoder and the ReferenceNet extract detailed features from the source face, which are then merged into the Stable Diffusion main UNet via spatial-attention. The face encoder (Deng et al. 2019) extracts high-level features from the source face image with querying transformer, which is then injected into both the ReferenceNet and the main UNet via cross-attention. Four composable face conditions are fed into four expert guiders and fused with fusion net in the latent space. Temporal-attention aims to improve the temporal consistency across frames. After iterative denoising, the output of the main UNet is decoded into the final animated video by a VAE decoder.
Identity-Preserving Generation
@@
The impressive generative abilities of foundation models have attracted recent research endeavors investigating their personalized generation potential. With popular pretrained T2I and T2V diffusion models like Stable Diffusion (Rombach et al. 2022) and AnimateDiff (Guo et al. 2024), several ID-Preserving methods are proposed with promising results. IP-Adapter (Ye et al. 2023) and InstantID (Wang et al. 2024a) could maintain high face fidelity with effective adaptable plugins. But, it’s hard to accurately control expression and pose without rich motion guidance, which makes video generation difficult. Animate Anyone (Hu 2024) intends to use spatial-attention to extract identity from source face image and pose sequence to offer motion information, but this lacks appearance alignment which results in inaccurate identity. Champ (Zhu et al. 2024) aims to introduce multiple pose conditions from 3D human parametric model for accurately capturing both pose and shape variations. However, the conditions of the champ are not sufficiently disentangled, leading to redundant information. In this paper, we aim to first decompose all facial conditions with their own unique usage and to make full use of the prior knowledge of Stable Diffusion for face swapping.
Method
@@
Preliminary: Latent Diffusion Models
Diffusion models generate a realistic image from a standard Gaussian distribution by reversing a recurrent noising process (Ho, Jain, and Abbeel 2020). However, pure diffusion process in pixel-space is really memory-consuming. Latent Diffusion Model (Rombach et al. 2022) makes the forward and reverse process in the latent space to reduce computation with an autoencoder. It first encodes the desired image into lower-dimension: $\mathbf{z}=\mathcal{E}(x)$. Then, LDM applies diffusion in the latent space. The forward process gradually alters to Gaussian distribution from the data $\mathbf{z}_{0}\sim q(\mathbf{z}_{0})$, which could be concluded as $q(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{t-1},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ is a predefined mean schedule. In addition, the reverse process is as follows:
@@
where $\mu_{\theta}$ and $\sigma_{\theta}$ are parameterized with neural network. A common training strategy that uses noise approximation model $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ performs better than other forms from experiments in previous work (Nichol and Dhariwal 2021). The training objective could be represented by a simplified L2 loss between ground truth noise and predicted noise:
$\mathcal{L}_{ldm}=\mathbb{E}_{\theta}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,c\right)\right\|^{2}\right],$ (2)
where $c$ is the designed facial conditions, and $t$ represents the timestep of current noisy image. At last, we use the decoder to reconstruct image from the latent output: $\hat{x}=\mathcal{D}(\mathbf{z}_{0})$.
Composable Face Conditions
The goal of face swapping is to preserve the identity of source face image and spatially align well with the target face image. Given a source face image $I_{src}$ and a sequence of target face video frames $I_{tgt}^{1:N}$, we first adopt a 3D facial reconstruction method to obtain 3D face model, $F_{src}$ and $F^{1:N}_{tgt}$, containing different geometry parameters like identity $\alpha$, pose $\beta$, expression $\theta$ and albedo $\gamma$. We intend to combine the identity of source face and motion of target face together. Figure 1 illustrates the attribute information contained in each condition and its corresponding source. The key of the composable face condition in our work is to design more accurate and independent condition for flexible and accurate controlling. To make conditions disentangled with each other, we decompose faces into four conditions: background, expression, shape, and illumination. The details are as follows:
Shape-Aware Pose Condition.
We use the 3DDFA-V3 (Wang et al. 2024b) to estimate facial parameters from each frame of target face video and source face image. Then, we render the 2D normal map from 3D face for pose and shape information. However, apart from the necessary pose and expression details, these images also encapsulate identity-specific facial geometry from the target individuals. The identity information essential for our purpose should be extracted from the source image. To address this mismatching issue, we utilize a facial shape alignment strategy which combines the identity information from the source image with the pose and expression information from the driving frames to generate the refined 2D facial conditions:
@@
Background-Preservation Condition.
We only need to change facial area in face swapping task. So, the background information could be directly sent to the model as a conditional inpainting task. The key of designing background condition is how to generate face mask. Most face swapping methods use the original facial mask of target face image, but this will ignore the face shape differences during inference. Face Adapter (Han et al. 2024) uses dilated facial area to train an area predictor for possible inpainting mask esstimation. Nevertheless, finding an optimal dilation size is crucial, as a face mask with small dilation also introduces useless shape information of the target face image, whereas a face mask with large dilation may contain face occlusion that makes model training more difficult. Here, we first use a pretrained face parser (Yu et al. 2018) to predict facial area $M_{tgt}$ and occlusion area $M_{occ}$. To avoid learning facial shape of target face during training, we apply random shift on the facial area $M_{tgt}$ to get $M_{tgt}^{shift}$. To address the shape misalignment problem in the inference phase, we replace $M_{tgt}^{shift}$ with $M_{tgt}^{normal}$ from the area of $I_{normal}$, containing shape of source face and position of target face. The swapping area $M_{swap}$ could be described as follow:
@@
Expression-Related Condition.
Facial expression is essential for face swapping. It’s hard to combine identity of source face image and facial expression of target face image. Some methods (Zhao et al. 2023; Han et al. 2024) use 3D face reconstruction to get 3D-aware keypoints by replacing the shape with source face image. However, the 3D-aware keypoints miss the detailed information of the eyes and its precision is suboptimal. In this work, we adopt 2D landmark of target face image, and transform points to segmentation for a simple and straight representation. We use several sets of boundary landmark to generate expression segmentation including areas of eyebrows, eyes, eyeballs and mouth:
@@
Identity-Erased Illumination Condition.
Environmental lighting is also a necessary condition to make generated face frames well aligned with target face frames. Few methods pay attention to illumination in face swapping task. With the 3D facial prior in previous section, we could use the UV texture map of target face frames to provide illumination information. UV texture map contains irrelevant identity of target face image, while we only need to capture lighting from it. So, we add blur on the rendered UV texture map to destroy target face identity as:
@@
Now, we have completed four composable face conditions with their unique usage by leveraging the 3D facial prior of both source face image and target face frames. All conditions are spatially mapped into image-level motion sequences. Inspired by the previous work (Zhu et al. 2024), we use Mixture-of-Guiders for lightweight guidance instead of several ControlNets. All guiders are implemented as a set of convolution and activation layers, aiming to extract local spatial information. This architecture facilitates condition adaptability while promoting parameter efficiency. We also add a self-attention module to learn the unique information from each disentangled face condition. This could capture more global information from each guidance. Even though all conditions are well designed for its distinct character and spatially aligned with generated face, there are still some mismatching which could not be solved due to limitation of 3D reconstruction and compression when passing to guiders. The conditions from some works are not disentangled from each other, which contains too much redundant information, and they directly fuse these guidance condition through summation. This would cause the network only learn useful information from some detailed condition and drop some of repeated condition for easier learning. We adopt a fusion net to better fuse these four distinct face conditions. We first concat all face conditions after each guiders and use a self-attention mechanism to learn the spatial relation among these conditions. Both guiders and fusion net apply zero convolution at the last layer to maintain the prior knowledge of original Stable Diffusion and make the whole architecture more lightweight to train with motion guidance.
Appearance-Aware Controlling
Face Former.

### v2

# DynamicFace: High-Quality and Consistent Face Swapping for Image and Video using Composable 3D Facial Priors

@@ Abstract
Face swapping transfers the identity of a source face to a target face while retaining the attributes like expression, pose, hair, and background of the target face. Advanced face swapping methods have achieved attractive results. However, these methods often inadvertently transfer identity information from the target face, compromising expression-related details and accurate identity. We propose a novel method DynamicFace that leverages the power of diffusion models and plug-and-play adaptive attention layers for image and video face swapping. First, we introduce four fine-grained facial conditions using 3D facial priors. All conditions are designed to be disentangled from each other for precise and unique control. Then, we adopt Face Former and ReferenceNet for high-level and detailed identity injection. Through experiments on the FF++ dataset, we demonstrate that our method achieves state-of-the-art results in face swapping, showcasing superior image quality, identity preservation, and expression accuracy. Our framework seamlessly adapts to both image and video domains. Our code and results will be available on the project page: https://dynamic-face.github.io/.
{strip}[Uncaptioned image]
1 Introduction
@@
Recent works [17, 4, 24] have made great efforts to achieve good face swapping results. However, these methods often focus on inner facial texture but ignore shape and illumination. Later, some works [28, 31, 19] combine the expert knowledge from 3D face reconstruction to achieve better performance. But training model with both reconstruction loss and identity loss will face a trade-off problem between identity and reconstruction, which often leads to instability and unreliability. Balancing identity and reconstruction with only losses constraint is conflict. We should provide accurate information for controllable generation.
Recently, diffusion-based models [11, 25, 23, 10] have exhibited high customizability for various conditions and impressive ability to generate images with high resolution and complex scenes. Some works [8, 34, 14] try to utilize the diversity and powerful generation ability of diffusion model to achieve face swapping. However, these methods could not achieve video face swapping directly. And idealized face guidance should preserve the non-identity attributes of $I_{tgt}$, which include non-facial attributes (e.g., background and hair), facial posture (e.g., expression and pose) and facial color (e.g., lighting). Previous works didn’t pay much attention to the potentiality of 3D facial priors to simultaneously maintain identity from source face and complicated motion information from target face.
In this work, we propose a method named DynamicFace for face swapping by incorporating precise and disentangled facial conditions to the powerful Stable Diffusion model [23] and apply temporal attention layer in AnimateDiff [7] to make DynamicFace also available in video domain. We introduce a new multiple conditioning mechanism to address the limitations of previous methods. Specifically, 1) to accurately control expression, pose, and other motion, we introduce four fine-grained facial conditions: background, shape-aware normal map, expression-related landmark, and identity-removed UV texture map. These four conditions fully encapsulate the desired information in the target face and are well disentangled from each other. 2) to avoid introducing unexpected identity information from the target face, we propose a alignment strategy to form spatial-aligned conditions using 3D facial priors. This involves extracting pose and expression parameters from the target face video and shape parameters from the source image using the prior of 3D face reconstruction model 3DDFA-V3 [29], and then rendering the 2D face normal map and the UV texture map with these facial parameters. In summary, our main contributions are as follows:
@@
  • We propose the Fusion TV Optimizer (FusionTVO), a plug-and-play inference module that innovatively combines position-weighted latent fusion and temporal TV regularization to eliminate segment flicker artifacts.
2 Related Work
@@ 2.1 Face Swapping
Early methods directly warp the source face according to the target facial landmarks, thus failing to address the large differences in posture and expression. 3DMM-based methods swap faces by 3D-fitting and re-rendering. However, these methods often cannot handle lighting differences on face and suffer from poor fidelity. Later, GAN-based methods improve the fidelity of the generated faces. Deepfakes [4] transfers the target attributes to the source face by an encoder-decoder structure while being constrained by two specific identities. FSGAN [21] animates the source face with target facial landmarks and fuses it with the target background. SimSwap [3] introduces a feature matching loss to help preserve the target attributes. These methods require a facial mask for blending, which limits face shape variation.
@@
Diffusion models have shown remarkable performance in various generative tasks, emerging as a powerful alternative to traditional GANs for image and video synthesis. The innovative approach of these models lies in their ability to gradually refine an image by reversing a learned noising process, which has led to significant advancements in image quality and diversity. In the context of face swapping, diffusion models have introduced a new paradigm for identity manipulation in face images. DiffSwap [34] proposes a diffusion-based pipeline in latent space for face swapping by integrating 3D-aware masked diffusion and midpoint estimation during the reverse process. DiffFace [14] introduces an ID conditional DDPM. However, pure pixel-wise diffusion without pretraining requires more data, and extra forward steps with gradients make training more memory-consuming. These both make training process difficult. Face Adapter [8] only utilizes an efficient adapter to introduce identity information and preserve attributes of the target face image, but it is difficult to achieve temporal consistency when balancing the trade-off between identity and motion. In addition, these diffusion-based methods could not easily transfer to the video domain.
@@
Figure 3: The overview of the proposed method. A VAE encoder and the ReferenceNet extract detailed features from the source face, which are then merged into the Stable Diffusion main UNet via spatial-attention. The face encoder [5] extracts high-level features from the source face image with querying transformer, which is then injected into both the ReferenceNet and the main UNet via cross-attention. Four composable face conditions are fed into four expert guiders and fused with fusion net in the latent space. Temporal-attention aims to improve the temporal consistency across frames. After iterative denoising, the output of the main UNet is decoded into the final animated video by a VAE decoder.
2.2 Identity-Preserving Generation
@@
The impressive generative abilities of foundation models have attracted recent research endeavors investigating their personalized generation potential. With popular pretrained T2I and T2V diffusion models like Stable Diffusion [23] and AnimateDiff [7], several ID-Preserving methods are proposed with promising results. IP-Adapter [32] and InstantID [27] could maintain high face fidelity with effective adaptable plugins. But, it’s hard to accurately control expression and pose without rich motion guidance, which makes video generation difficult. Animate Anyone [12] intends to use spatial-attention to extract identity from source face image and pose sequence to offer motion information, but this lacks appearance alignment which results in inaccurate identity. Champ [35] aims to introduce multiple pose conditions from 3D human parametric model for accurately capturing both pose and shape variations. However, the conditions of the champ are not sufficiently disentangled, leading to redundant information. In this paper, we aim to first decompose all facial conditions with their own unique usage and to make full use of the prior knowledge of Stable Diffusion for face swapping.
3 Method
@@
3.1 Preliminary: Latent Diffusion Models
Diffusion models generate a realistic image from a standard Gaussian distribution by reversing a recurrent noising process [11]. However, pure diffusion process in pixel-space is really memory-consuming. Latent Diffusion Model [23] makes the forward and reverse process in the latent space to reduce computation with an autoencoder. It first encodes the desired image into lower-dimension: $\mathbf{z}=\mathcal{E}(x)$. Then, LDM applies diffusion in the latent space. The forward process gradually alters to Gaussian distribution from the data $\mathbf{z}_{0}\sim q(\mathbf{z}_{0})$, which could be concluded as $q(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{t-1},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ is a predefined mean schedule. In addition, the reverse process is as follows:
@@ 3.1 Preliminary: Latent Diffusion Models
where $\mu_{\theta}$ and $\sigma_{\theta}$ are parameterized with neural network. A common training strategy that uses noise approximation model $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ performs better than other forms from experiments in previous work [20]. The training objective could be represented by a simplified L2 loss between ground truth noise and predicted noise:
$\mathcal{L}_{ldm}=\mathbb{E}_{\theta}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\mathbf{z}_{t},t,c\right)\right\|^{2}\right],$ (2)
where $c$ is facial conditions, and $t$ represents the timestep of current noisy image. At last, we use the decoder to reconstruct image from the latent output: $\hat{x}=\mathcal{D}(\mathbf{z}_{0})$.
3.2 Composable Face Conditions
The goal of face swapping is to preserve the identity of source face image and spatially align well with the target face image. Given a source face image $I_{src}$ and a sequence of target face video frames $I_{tgt}^{1:N}$, we first adopt a 3D facial reconstruction method to obtain 3D face model, $F_{src}$ and $F^{1:N}_{tgt}$, containing different geometry parameters like identity $\alpha$, pose $\beta$, expression $\theta$ and albedo $\gamma$. We intend to combine the identity of source face and motion of target face together. Figure 2 illustrates the attribute information contained in each condition and its corresponding source. The key of the composable face condition in our work is to design more accurate and independent condition for flexible and accurate controlling. To make conditions disentangled with each other, we decompose faces into four conditions: background, expression, shape, and illumination. The details are as follows: Shape-Aware Pose Condition. We use the 3DDFA-V3 [29] to estimate facial parameters from each frame of target face video and source face image. Then, we render the 2D normal map from 3D face for pose and shape information. However, apart from the necessary pose and expression details, these images also encapsulate identity-specific facial geometry from the target individuals. The identity information essential for our purpose should be extracted from the source image. To address this mismatching issue, we utilize a facial shape alignment strategy which combines the identity information from the source image with the pose and expression information from the driving frames to generate the refined 2D facial conditions:
@@ 3.2 Composable Face Conditions
Background-Preservation Condition. We only need to change facial area in face swapping task. So, the background information could be directly sent to the model as a conditional inpainting task. The key of designing background condition is how to generate face mask. Most face swapping methods use the original facial mask of target face image, but this will ignore the face shape differences during inference. Face Adapter [8] uses dilated facial area to train an area predictor for possible inpainting mask esstimation. Nevertheless, finding an optimal dilation size is crucial, as a face mask with small dilation also introduces useless shape information of the target face image, whereas a face mask with large dilation may contain face occlusion that makes model training more difficult. Here, we first use a pretrained face parser [33] to predict facial area $M_{tgt}$ and occlusion area $M_{occ}$. To avoid learning facial shape of target face during training, we apply random shift on the facial area $M_{tgt}$ to get $M_{tgt}^{shift}$. To address the shape misalignment problem in the inference phase, we replace $M_{tgt}^{shift}$ with $M_{tgt}^{normal}$ from the area of $I_{normal}$, containing shape of source face and position of target face. The swapping area $M_{swap}$ could be described as follow:
@@
Expression-Related Condition. Facial expression is essential for face swapping. It’s hard to combine identity of source face image and facial expression of target face image. Some methods [34, 8] use 3D face reconstruction to get 3D-aware keypoints by replacing the shape with source face image. However, the 3D-aware keypoints miss the detailed information of the eyes and its precision is suboptimal. In this work, we adopt 2D landmark of target face image, and transform points to segmentation for a simple and straight representation. We use several sets of boundary landmark to generate expression segmentation including areas of eyebrows, eyes, eyeballs and mouth:
@@
Identity-Erased Illumination Condition. Environmental lighting is also a necessary condition to make generated face frames well aligned with target face frames. Few methods pay attention to illumination in face swapping task. With the 3D facial prior in previous section, we could use the UV texture map of target face frames to provide illumination information. UV texture map contains irrelevant identity of target face image, while we only need to capture lighting from it. So, we add blur on the rendered UV texture map to destroy target face identity as:
@@ 3.3 Mixture-of-Guiders
Now, we have completed four composable face conditions with their unique usage by leveraging the 3D facial prior of both source face image and target face frames. All conditions are spatially mapped into image-level motion sequences. Inspired by the previous work [35], we use Mixture-of-Guiders for lightweight guidance instead of several ControlNets. All guiders are implemented as a set of convolution and activation layers, aiming to extract local spatial information. This architecture facilitates condition adaptability while promoting parameter efficiency. We also add a self-attention module to learn the unique information from each disentangled face condition. This could capture more global information from each guidance. Even though all conditions are well designed for its distinct character and spatially aligned with generated face, there are still some mismatching which could not be solved due to limitation of 3D reconstruction and compression when passing to guiders. The conditions from some works are not disentangled from each other, which contains too much redundant information, and they directly fuse these guidance condition through summation. This would cause the network only learn useful information from some detailed condition and drop some of repeated condition for easier learning. We adopt a fusion net to better fuse these four distinct face conditions. We first concat all face conditions after each guiders and use a self-attention mechanism to learn the spatial relation among these conditions. Both guiders and fusion net apply zero convolution at the last layer to maintain the prior knowledge of original Stable Diffusion and make the whole architecture more lightweight to train with motion guidance.
3.4 Appearance-Aware Controlling